Use Cramer's rule to find the solution of the following systems of linear equation in terms of the parameter K.

a) 5x - ky = 6. b) 2x - 3y = k

- 2x + 2ky = -3 X + 2y = - 2


1
Expert's answer
2021-04-29T17:43:48-0400

a)

"\begin{cases}
5x-ky = 6
\\
-2x+2ky = -3
\end{cases}
\\
\Delta =\begin{vmatrix}
5 & -k \\
-2 & 2k 
\end{vmatrix}
=10k -2k = 8k \\
\Delta_1 = 
\begin{vmatrix}
6 &-k\\
3&2
\end{vmatrix} = 12k-3k = 9k\\
\Delta_2 = 
\begin{vmatrix}
5&6\\
-2&-3
\end{vmatrix} = -15+12 = -3
\\
x = \cfrac{\Delta_1}{\Delta} = \cfrac{9k}{8k} = \cfrac{9}{8}\\
y = \cfrac{\Delta_2}{\Delta} =\cfrac{-3}{8k}"

b)

"\begin{cases}
2x - 3y = k\\
x+2y = -2
\end{cases}\\
\Delta = 
\begin{vmatrix}
2 & -3 \\ 
1 & 2
\end{vmatrix} = 7\\
\Delta_1 = 
\begin{vmatrix}
k&-3\\
-2&2
\end{vmatrix} =2k-6\\
\Delta_2= 
\begin{vmatrix}
2&k\\
1&-2
\end{vmatrix} = -4-k\\
x = \cfrac{\Delta_1}{\Delta} =\cfrac{2k-6}{7}\\
y = \cfrac{\Delta_2}{\Delta} = -\cfrac{4+k}{7}"